OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6.

ganeshie8 (ganeshie8):

equation of a line passing through the point (x1, y1) with a slope of m : y-y1 = m(x-x1) in this problem, slope m = ?